Learned Corporation has provided the following information: Cost per Unit Cost per Period Direct materials $ 5.40 Direct labor $ 3.80 Variable manufacturing overhead $ 1.40 Fixed manufacturing overhead $ 38,250 Sales commissions $ 0.80 Variable administrative expense $ 0.70 Fixed selling and administrative expense $ 12,750 


Required:

 a. For financial reporting purposes, what is the total amount of product costs incurred to make 8,500 units?

 b. For financial reporting purposes, what is the total amount of period costs incurred to sell 8,500 units?

 c. If the selling price is $22.90 per unit, what is the contribution margin per unit sold? (Round your answer to 2 decimal places.)

 d. If 9,500 units are produced, what is the total amount of direct manufacturing cost incurred?

 e. If 9,500 units are produced, what is the total amount of indirect manufacturing costs incurred?


a. Total product (manufacturing) cost

b. Total period (nonmanufacturing) cost

c. Contribution margin per unit

d. Total direct manufacturing cost

e. Total indirect manufacturing cost
Answer #1
a
Direct materials 45900 =8500*5.4
Direct labor 32300 =8500*3.8
Variable manufacturing overhead 11900 =8500*1.4
Fixed manufacturing overhead 38250
Total product (manufacturing) cost 128350
b
Sales commissions 6800 =8500*0.8
Variable administrative expense 5950 =8500*0.7
Fixed selling and administrative expense 12750
Total period (nonmanufacturing) cost 25500
c
Selling price 22.90
Less: Variable costs per unit
Direct materials 5.40
Direct labor 3.80
Variable manufacturing overhead 1.40
Sales commissions 0.80
Variable administrative expense 0.70
Total Variable costs per unit 12.10
Contribution margin per unit 10.80
d
Direct materials 51300 =9500*5.4
Direct labor 36100 =9500*3.8
Total direct manufacturing cost 87400
e
Variable manufacturing overhead 13300 =9500*1.4
Fixed manufacturing overhead 38250
Total indirect manufacturing cost 51550